how do I simplify the problem:
ln(x+2)+ln3=ln39

Answer by AdolphousC(70)

ln(x+2)+ln3=ln39
You should be able to use the product property to rewrite the problem as follows
ln 3(x+2) = ln 39
the if the ln 3(x+2) = ln 39 then 3(x+2) = 39


now solve like normal
3x + 6 = 39
3x = 33
x = 11